OpenStudy (anonymous):

Rose and Dennis each open a savings account at the same time. Rose invests $2,600 in an account yielding 4.1% simple interest, and Dennis invests $2200 in an account yielding 5.7% simple interest. After nine years, who has the greater total amount of money, and how much greater is it?

OpenStudy (anonymous):

Rose: A=P(1+RT) A=$2600(1+0.041*9)=$3559.40 Dennis: A=$2200(1+0.057*9)=$3328.60 Rose-Dennis: $3559.40-$3328.60=$230.80 Rose would have $3559.40 and Dennis would have $3328.60. Rose would have $230.80 more than dennis

OpenStudy (anonymous):

Thank you so much , i suck with this interest stuff .. if i have another question could u help me right away ??

OpenStudy (anonymous):

if i can answer it sure

OpenStudy (anonymous):

ok thank u :}

OpenStudy (anonymous):

Mary and Lewis each open a savings account at the same time. Mary invests $3,700 in an account yielding 3.2% simple interest, and Lewis invests $3,000 in an account yielding 5.9% simple interest. After fifteen years, who has the greater total amount of money, and how much greater is it?

OpenStudy (anonymous):

so basically same thing using the formula A=P(1+RT) P=principal R=rate T=time For mary: A=$3700(1+0.032*15) A=$5476 For Lewis:A=$3000(1+0.059*15) A=$5655 So who has greater take the person who gets more subtract who has less in this case lewis-mary = $5655-5476=$179 So Lewis has a greather amount of money by $179

OpenStudy (anonymous):

3.2%=3.2/100=0.032 5.7%=5.7/100=0.057

OpenStudy (anonymous):

omg thank u soo much u have been the biggest help !! :} i wish i could give u something lol or a hug haha !! well thank u :)

OpenStudy (anonymous):

your welcome.

OpenStudy (anonymous):

any other question?

OpenStudy (anonymous):

no not as of now but thank u !!:) if i do ill make sure ill ask u :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok soo im in compound interest and sooo confused lol

OpenStudy (anonymous):

ok what are the questions?

OpenStudy (anonymous):

You have $12,000 to invest and want to keep your money invested for 8 years. You are considering the following investment options. Choose the investment option that will earn you the most money.

OpenStudy (anonymous):

options?

OpenStudy (anonymous):

a. 3.99% compounded monthly b. 4% compounded quarterly c. 4.175% compounded annually d. 4.2% simple interest

OpenStudy (anonymous):

ok for compound interest use the formula: V = P(1+r/n)^nt V = the future value of the investment P = the principal investment amount r = the annual interest rate n = the number of times that interest is compounded per year t = the number of years the money is invested for

OpenStudy (anonymous):

so for part a: it's V=$12,000(1+0.0399/12)^12*8 =$12,000(1.003325)^96=$12,000(1.375298094)=$16,503.58 n=12 because number of times a year is 12 months in a year Part b: V=$12,000(1+0.04/4)^4*8 V=$12,000(1.01)^32 V=$12,000(1.374940679)=$16,499.28814 or $16,499.29 n=4 because number of times a year is 4 quarters in a year Part C: V=$12,000(1+0.04175/1)^1*8 V=$12,000(1.387100962)=$16,645.21155 or $16,645.21 n=1 only once per year Part d: A=P(1+rt) A=$12,000(1+0.042*8) A=$12,000(1.336)=$16,032 So option c would earn you the most money.

OpenStudy (anonymous):

thank you again soo much !! you're awesome !!:)

OpenStudy (anonymous):